# Improving Research Idea Generation Through Data: An Empirical Investigation in Social Science

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have shown promise in generating novel research ideas. However, these ideas often face challenges related to feasibility and expected effectiveness. This paper explores how augmenting LLMs with relevant data during the idea generation process can enhance the quality of generated ideas. We introduce two ways of incorporating data: (1) providing metadata during the idea generation stage to guide LLMs toward feasible directions, and (2) adding automatic validation during the idea selection stage to assess the empirical plausibility of hypotheses within ideas. We conduct experiments in the social science domain, specifically with climate negotiation topics, and find that metadata improves the feasibility of generated ideas by 20%, while automatic validation improves the overall quality of selected ideas by 7%. A human study shows that LLM-generated ideas, along with their related data and validation processes, inspire researchers to propose research ideas with higher quality. Our work highlights the potential of data-driven research idea generation, and underscores the practical utility of LLM-assisted ideation in real-world academic settings.
Improving Research Idea Generation Through Data: An Empirical Investigation in Social Science
Xiao Liu1, Xinyi Dong2, Xinyang Gao3, Yansong Feng1111Corresponding authors.and Xun Pang4,5,6111Corresponding authors. 1Wangxuan Institute of Computer Technology  2Yuanpei College  3School of Government 4School of International Studies  5Institute for Carbon Neutrality 6Analytics Lab for Global Risk Politics Peking University {lxlisa,xinyang0614,fengyansong,xpang}@pku.edu.cn dongxy@stu.pku.edu.cn
@@ 1 Introduction
Figure 1: Overview of how we incorporate data into the research idea generation process.
@@
Recent advances in large language models (LLMs) have demonstrated their potential to generate domain-specific research ideas, with some studies suggesting that these ideas can exhibit greater novelty than those proposed by human experts (Si et al., 2024; Yamada et al., 2025). However, many LLM-generated ideas suffer from practical limitations: they may be infeasible to implement, lack suitable datasets for validation, or have uncertain effectiveness. For instance, an LLM might propose investigating “the impact of diplomats’ childhood environmental experiences on their bargaining positions in UN climate negotiations”, which is an interesting idea but lacks available data for empirical analysis.
@@
Intuitively, if LLMs are provided with relevant datasets, they could be better equipped to generate empirically grounded research ideas: those that are not only novel but also feasible for experimentation. Just as human researchers navigate trade-offs between theoretical ambition and empirical tractability when developing research ideas, LLMs could benefit from this balancing act when data is available. For example, if the LLM is aware of the existence of records on climate conference attendance, it might propose a more feasible study, like “how the professional backgrounds of diplomats influence their countries’ emission reduction commitment ambitions.”
@@
In this paper, we investigate whether augmenting LLMs with data during the research idea generation process can enhance the quality of generated ideas. Data can not only help LLMs to generate more feasible ideas, but also enable preliminary validation of hypotheses within ideas. With access to relevant datasets, LLMs can write code to analyze the data and perform reasoning to assess whether the hypotheses are supported by the available evidence. Although this validation is preliminary and does not guarantee sound conclusions, it provides valuable signals regarding whether the ideas are likely to be effective.
@@
As shown in Figure 1, the standard framework for LLM ideation consists of three stages: literature search, idea generation, and idea selection. Models first search related literature for a given topic, then generate ideas based on the retrieved literature, and finally rank and select the top ideas as the output. We enhance this framework by incorporating data at two key stages: (1) During idea generation, we provide metadata, such as dataset descriptions, to guide models toward feasible research directions; and (2) during idea selection, we integrate automatic validation to account for the empirical plausibility of the proposed hypotheses within ideas.
@@
We conduct experiments in the domain of social science, focusing specifically on topics related to climate negotiations. To support the experiments, we first collect and gather relevant datasets into a unified ClimateDataBank. To evaluate the impact of metadata, we compare LLM-generated ideas with and without access to the metadata of ClimateDataBank, and observe that incorporating metadata improves the feasibility by 20% and the expected effectiveness by 18% in human evaluation. Additionally, we find that automatic validation improves the accuracy of idea ranking by an average of 8%, and ideas selected with validation are rated 7% higher in human evaluation compared to those selected without validation.
@@
Our contributions are as follows: (1) We propose two ways of integrating data into research idea generation: adding metadata in idea generation and adding automatic validation in idea selection. (2) Our experiments demonstrate that metadata and automatic validation improve the quality of generated ideas, particularly in feasibility and expected effectiveness. (3) The human study reveals that LLM-generated ideas can inspire researchers to propose higher-quality ideas. (4) We construct the ClimateDataBank to support future work in data-driven research idea generation.
@@ 2 Data Collection
We first collect data related to climate negotiations, constructing the ClimateDataBank to facilitate the following experiments. Our process begins with a comprehensive literature review, identifying important and commonly used datasets. We then collect datasets of common variables from World Bank Open Data***https://data.worldbank.org/, and other datasets from their original sources.
@@
Figure 2: An example of metadata provided during idea generation, alongside a generated research idea.
3 Can Metadata Benefit Idea Generation?
@@ 3.1 Social Science Idea Generation
A typical social science research idea consists of three components: a research question $r_{q}$, a theory $t_{h}$, and several hypotheses $\boldsymbol{h}$ (Powner, 2014; King et al., 1994). As illustrated in the example of Figure 2 (right), the research question guides the study by identifying the central issue to be explored. The theory speculates on the answer to the research question and explains why the proposed answer is reasonable. The hypotheses identify observable implications of the theory, i.e., things we would observe if the theory is correct.
@@ 3.2 Incorporating Metadata into Idea Generation
Figure 2 (left) shows how we incorporate metadata, which is concise dataset descriptions, into the idea generation prompt along with the topic and related literature. Each metadata entry summarizes a dataset in ClimateDataBank with one or two sentences, including information like the meaning of key variables, temporal coverage, and spatial scope. The prompt informs LLMs that here are existing data related to this topic, without strict restriction on using the provided data. This ensures that models can balance theoretical creativity with empirical feasibility by themselves.
@@
Figure 3: Automatic evaluation results of ideas generated with (w.) and without metadata. A tabular version is in Appendix Table 10.
@@ 3.3 Experimental Setup
We generate 10 climate negotiation-related research topics using GPT-4o (Hurst et al., 2024), and manually verify them to ensure their quality. The created topics are in Appendix Table 8.
@@ Methods
We experiment with three prevalent research idea generation methods: AI-Researcher (Si et al., 2024), GPT-Researcher (Elovic, 2023), and Chain-of-Ideas (Li et al., 2024a). Each method first retrieves relevant literature and then generates ideas, with detailed descriptions in Appendix C. We preserve each method’s original design while adding metadata to the idea generation prompt. For each research topic, we generate 50 ideas per method, then select the top 5 for evaluation using a unified idea selection module.
@@
As Si et al. (2024) demonstrate that LLMs better assess ideas in pairwise ranking than rating, we conduct a Swiss tournament for idea selection. Over 5 rounds, ideas are paired by similar accumulated scores, with LLMs ranking each pair using the criteria below.
@@
We assess idea quality using four criteria motivated by previous works (Si et al., 2024; Yang et al., 2024b): significance, novelty, feasibility, and expected effectiveness (abbreviated as exp. effectiveness). These criteria are used in both idea selection and evaluation, with detailed definitions in Appendix D. For automatic evaluation, we conduct tournament ranking and compute ELO scores following Idea Arena (Li et al., 2024a).
@@
We use GPT-4o (gpt-4o-2024-08-06) for idea generation and selection, and use Gemini-1.5-Pro (Team et al., 2024) (gemini-1.5-pro-002) and Claude-3.5-Sonnet (Anthropic, 2024) (claude-3-5-sonnet-20241022) as judge models.These model versions are used throughout the paper. More implementation details, prompts, and human evaluation details are in Appendices E, H, and I, respectively.
@@ 3.4 Results
Figure 3 shows that metadata improves average ratings across all methods, suggesting that incorporating metadata enhances the overall quality of generated research ideas. Expected effectiveness consistently benefits from metadata, with feasibility and significance also improving in most cases, demonstrating metadata’s role in generating more empirically grounded and impactful ideas. However, novelty declines for AI-Researcher and Chain-of-Ideas when evaluated by Claude, indicating that data-aware generation may limit the generation of highly unconventional ideas.
@@
w. Metadata Tie w/o Metadata
Significance 38.8 22.4 38.8
Novelty 42.6 14.0 43.4
Feasibility 46.5 27.1 26.4
Exp. Effectiveness 51.2 16.3 32.5
Overall 43.4 14.7 41.9
[…]
We perform human evaluation on GPT-Researcher’s output, as this method achieves high rankings in the automatic evaluation. We recruit human annotators at the graduate level or above, with academic backgrounds in social science. For the 50 idea pairs (5 ideas per topic across 10 topics) generated by GPT-Researcher with and without metadata, each pair is annotated by at least two participants. In addition to the four evaluation criteria, we also ask annotators to assess the overall quality of research ideas.
As shown in Table 1, the human evaluation results align with the automatic evaluation trends. The integration of metadata leads to substantial improvements in feasibility (20%) and expected effectiveness (18%), though novelty experiences a modest decrease. The overall assessment score increases by 1.5%, suggesting that while metadata strengthens specific quality dimensions, the holistic assessment of research ideas involves balancing multiple criteria. Appendix G demonstrates an example where incorporating metadata improves the feasibility of generated idea, thus improving the overall quality of the idea.
[…]
Diverse 20 100 78.0
Climate 8 18 72.2
(a) Accuracy of the hypothesis validation results.
[…]
Mostly Validate the Hypothesis 50
Partially Validate the Hypothesis 40
Does not Help Validating 10
(b) Human evaluation results of the validation processes.
30.0 63.3 30.0 33.3
[…]
4 Can Automatic Validation Benefit Idea Selection?
This section investigates the role of automatic validation in idea selection through: (1) a pilot study assessing the hypothesis validation capabilities of LLMs, (2) a reference-based evaluation of idea selection performance, and (3) a human evaluation comparing ideas selected with and without the validation process.
4.1 Pilot Study: Can LLMs conduct preliminary hypothesis validation?
Experimental Setup
We extract 18 hypotheses from the 8 domain-specific papers collected in §2, of which 10 are supported and 8 are refuted. Hypotheses with insignificant or mixed evidence are excluded. To expand the experimental scope, we sample 50 hypotheses from DiscoveryBench (Majumder et al., 2024b), drawn from 20 papers across diverse fields like humanities, sociology, and economics. Since all these hypotheses are supported in the original papers, we create 50 negative hypotheses by modifying their variables or relations to balance the evaluation dataset.
We experiment with GPT-4o using the code interpreter assistant, a built-in tool available in GPT models. It achieves superior performance in quantitative reasoning with data (Liu et al., 2024b), while more advanced methods can also be employed in the future. We input the hypotheses along with their corresponding data into the model. The model engages in multi-turn interactions to write and run Python code in a sandbox environment, to validate whether the hypotheses are supported.
Results
Figure 4: An example of incorporating automatic validation into idea selection. We check the feasibility of hypotheses in ideas, conduct automatic validation of feasible hypotheses, and provide the ideas together with the summarized validation processes to the judge model.
@@ 4.2 Incorporating Automatic Validation into Idea Selection
Figure 4 illustrates how we conduct automatic validation of generated ideas and incorporate the validation results into the idea selection process.
@@
If the idea is deemed testable, the corresponding datasets are then provided to the LLM for validation. Similar to the pilot study, the model generates a validation trace consisting of code snippets, intermediate execution outputs, and textual interpretations.
@@
We use GPT-4o for both the feasibility check and validation process summarization, and GPT-4o with the code interpreter assistant for hypothesis validation. Implementation details and prompts are in Appendices E and H. A detailed case of the automatic validation process is in Appendix G.
@@ 4.3 Reference-based Automatic Evaluation
[…]
We evaluate the impact of validation on idea selection performance, following the setup of ResearchBench (Liu et al., 2025). We prompt LLMs to perform pairwise ranking between ground-truth ideas (extracted from academic papers) and LLM-generated ideas, and compare ranking accuracy with and without access to validation processes.
@@
We use Gemini-1.5-Pro and Claude-3.5-Sonnet as judge LLMs. Results are shown in Table 3. For both models, incorporating validation leads to consistently higher average ranking accuracy compared to the setting without validation. Improvements are particularly notable in the feasibility and expected effectiveness dimensions, while a slight decrease is observed in the judgment of significance and novelty.

### v2

# arXiv:2505.21396v2[cs.CL] 28 Feb 2026

@@ Abstract
Recent advancements in large language models (LLMs) demonstrate strong potential for generating novel research ideas, yet such ideas often struggle with feasibility and effectiveness. In this paper, we investigate whether augmenting LLMs with relevant data during the ideation process can improve idea quality. Our framework integrates data at two stages: (1) incorporating metadata during idea generation to guide models toward more feasible concepts, and (2) introducing an automated preliminary validation step during idea selection to assess the empirical plausibility of hypotheses within ideas. We evaluate our approach in the social science domain, with a specific focus on climate negotiation topics. Expert evaluation shows that metadata improves the feasibility of generated ideas by 20%, while automated validation improves the overall quality of selected ideas by 7%. Beyond assessing the quality of LLM-generated ideas, we conduct a human study to examine whether these ideas, augmented with related data and preliminary validation, can inspire researchers in their own ideation. Participants report that the LLM-generated ideas and validation are highly useful, and the ideas they propose with such support are proven to be of higher quality than those proposed without assistance. Our findings highlight the potential of data-augmented research ideation and underscore the practical value of LLM-assisted ideation in real-world academic settings.
@@ 1 Introduction
Figure 1: Overview of the data-augmented LLM ideation framework. Compared to the standard ideation framework (center), our approach integrates metadata into the idea generation stage (top) and adds preliminary automatic validation to the idea selection stage (bottom).
@@
Recent advances in large language models (LLMs) have shown promise in generating domain-specific research ideas, with some studies suggesting that these ideas can exhibit greater novelty than those proposed by human experts [26, 32]. However, many LLM-generated ideas suffer from practical limitations: they may be infeasible to implement, lack suitable datasets for validation, or have uncertain effectiveness. For instance, an LLM might propose investigating “the impact of diplomats’ childhood environmental experiences on their bargaining positions in UN climate negotiations”, which is an interesting idea but lacks available data for empirical analysis.
@@
A potential reason is that current ideation methods mainly rely on literature, without guidance from empirical data [33, 15]. Intuitively, if LLMs are provided with relevant datasets, they could be better equipped to generate empirically grounded research ideas: those that are not only novel but also feasible for experimentation. Just as human researchers navigate trade-offs between theoretical ambition and empirical tractability when developing research ideas, LLMs could benefit from this balancing act when data is available. For example, if the LLM is aware of the existence of records on climate conference attendance, it might propose a more feasible study, like “how the professional backgrounds of diplomats influence their countries’ emission reduction commitment ambitions.”
@@
In this paper, we investigate whether augmenting LLMs with data during the ideation process can enhance the quality of generated ideas. Data can not only help LLMs to generate more feasible ideas, but also enable preliminary validation of hypotheses within ideas. With access to relevant datasets, LLMs can write code to analyze the data and perform reasoning to assess whether the hypotheses are supported by the available evidence. Although this validation is preliminary and does not guarantee sound conclusions, it provides valuable signals regarding whether the ideas are likely to be effective.
@@
As shown in the center of Figure 1, the standard framework for LLM ideation consists of three stages: literature search, idea generation, and idea selection. Models first search related literature for a given topic, then generate ideas based on the retrieved literature, and finally rank and select the top ideas as the output. We enhance this framework by incorporating data at two key stages: (1) During idea generation, we provide metadata, such as dataset descriptions, to guide models toward feasible research directions; and (2) during idea selection, we integrate automatic validation to account for the empirical plausibility of the proposed hypotheses within ideas.
@@
We conduct experiments in the domain of social science, focusing specifically on topics related to climate negotiations. Because social science research often deals with complex, context-dependent phenomena, access to high-quality data is essential for grounding theories in observable evidence and distinguishing between ideas that are merely interesting and those that are empirically testable. To support the experiments, we first collect and gather relevant datasets into a unified ClimateDataBank. We compare LLM-generated ideas with and without the metadata of ClimateDataBank, and observe that incorporating metadata improves the feasibility by 20% and the expected effectiveness by 18% in human evaluation. Furthermore, we compare ideas selected with and without access to the automatic validation process, and find that the overall quality of ideas selected with validation is rated 7% higher by human experts.
@@
Our contributions are as follows: (1) We propose two ways of integrating data into ideation: adding metadata in idea generation and adding automatic validation in idea selection. (2) Our experiments demonstrate that metadata and automatic validation improve the quality of generated ideas, particularly in feasibility and expected effectiveness. (3) Human study reveals that LLM-generated ideas can inspire researchers to propose higher-quality ideas, demonstrating the practical value of LLM ideation. (4) We construct the ClimateDataBank to support future work in data-driven ideation.
2 Related Work
Table 1: Comparison with existing works on research ideation and automated scientific discovery.
Work Primary Focus Data Integration Stage Help to Researchers
Si et al. [26] Generate research ideas Not explored
Baek et al. [2] Generate research ideas Not explored
Jansen et al. [12] End-to-end science discovery Experiment execution Not explored
Lu et al. [20] End-to-end science discovery Experiment execution Not explored
Ours Generate more feasible Idea generation Inspire researchers in their
and effective ideas and selection own ideation process
Research Idea Generation
There is growing interest in leveraging LLMs for research idea generation, either as a standalone task [26, 2] or as part of an end-to-end automated research pipeline [16, 20, 12]. The first category of work focuses on enhancing literature search and idea formulation, typically generating ideas grounded in prior work [29, 33]. The second category of work proposes more comprehensive frameworks that encompass later research stages, such as experiment design, execution, and paper writing.
Table 1 summarizes how our work differs from prior studies. The focus of our work aligns closely with the first category, but novelly incorporates data into the ideation process, aiming to generate more feasible and effective ideas. While our work also involves code generation and execution for hypothesis validation, this is not intended as rigorous experiment execution, but serves as a preliminary signal to support idea selection, which differs from the second category. Moreover, our work investigates how the generated ideas can support human researchers, which is rarely covered in previous research.
Hypothesis Generation
A related but distinct line of work focuses on hypothesis generation, where models generate hypotheses to explain phenomena given access to data, like inducing rules from observations [35, 25]. Studies in this area explore data-driven methods [21, 36] or integrate literature with data [17], but their goal is to uncover patterns in existing datasets, contrasting with our objective of generating high-quality research ideas.
@@ 3 Data Collection
We first collect data related to climate negotiations, constructing the ClimateDataBank to facilitate following experiments. Our process begins with a comprehensive literature review, identifying important and commonly used datasets. We collect common variables from World Bank Open Data [^1] https://data.worldbank.org/, and other datasets from their original sources.
@@
Given the cost of constructing the data bank and recruiting domain experts to conduct human evaluation, we restrict experiments to research topics related to climate negotiations. However, the framework we propose can be broadly applied to diverse quantitative social science studies, where researchers need to build a theory for a phenomenon and test hypotheses to support the theory.
4 Incorporating Metadata in Idea Generation
@@ 4.1 Social Science Idea Generation
A typical social science research idea consists of three components: a research question $r_{q}$, a theory $t_{h}$, and several hypotheses $\boldsymbol{h}$ [24, 14]. As illustrated in the top right example of Figure 1, the research question guides the study by identifying the central issue to be explored. The theory speculates on the answer to the research question and explains why the proposed answer is reasonable. The hypotheses identify observable implications of the theory, i.e., things we would observe if the theory is correct.
@@ 4.2 Incorporating Metadata into Idea Generation
Figure 1 (top) shows how we incorporate metadata, which is concise dataset descriptions, into the idea generation stage along with the topic and related literature. Each metadata entry summarizes a dataset in ClimateDataBank with one or two sentences, including information like the meaning of key variables, temporal coverage, and spatial scope. The prompt informs LLMs that here are existing data related to this topic, without strict restriction on using the provided data. This ensures that models can balance theoretical creativity with empirical feasibility by themselves.
@@
Table 2: ELO scores of ideas generated with (w.) and without metadata. Better results are in bold.
Gemini-1.5-Pro as the Judge
AI-Researcher 902 933 1047 951 958
938 931 1098 997 991
GPT-Researcher 1019 1000 1045 1015 1020
1073 1021 1183 1134 1103
Chain-of-Ideas 974 1025 822 915 934
1094 1091 805 988 995
Claude-3.5-Sonnet as the Judge
AI-Researcher 870 968 1060 881 945
855 859 1152 918 946
GPT-Researcher 931 972 1076 928 977
1000 903 1228 1085 1054
Chain-of-Ideas 1066 1118 768 1012 991
1278 1180 716 1176 1088
@@ 4.3 Experimental Setup
We generate 10 climate negotiation-related research topics using GPT-4o [11], and manually verify them to ensure their quality. The created topics are in Appendix Table 9.
@@ Methods
We experiment with three prevalent research idea generation methods: AI-Researcher [26], GPT-Researcher [8], and Chain-of-Ideas [15]. Each method first retrieves relevant literature and then generates ideas, with detailed descriptions in Appendix C. We preserve each method’s original design while adding metadata to the idea generation prompt. For each research topic, we generate 50 ideas per method, then select the top 5 for evaluation using a unified idea selection module.
@@
As Si et al. [26] demonstrate that LLMs better assess ideas in pairwise ranking than rating, we conduct a Swiss tournament for idea selection. Over 5 rounds, ideas are paired by similar accumulated scores, with LLMs ranking each pair using the criteria below.
@@
We assess idea quality using four criteria motivated by previous works [26, 34]: significance, novelty, feasibility, and expected effectiveness (abbreviated as exp. effectiveness). These criteria are used in both idea selection and evaluation, with definitions in Appendix D. For automatic evaluation, we conduct tournament ranking and compute ELO scores following Idea Arena [15].
@@
We use GPT-4o (gpt-4o-2024-08-06) for idea generation and selection, and use Gemini-1.5-Pro [27] (gemini-1.5-pro-002) and Claude-3.5-Sonnet [1] (claude-3-5-sonnet-20241022) as judge models. [^2] These model versions are used throughout the paper. More implementation details, prompts, and evaluation details are in Appendices E, H, and I.
@@ 4.4 Results
Table 2 shows that metadata improves average ratings across all methods, suggesting that incorporating metadata enhances the overall quality of generated research ideas. Expected effectiveness consistently benefits from metadata, with feasibility and significance also improving in most cases, demonstrating metadata’s role in generating more empirically grounded and impactful ideas. However, novelty declines for AI-Researcher and GPT-Researcher when evaluated by Claude, indicating that data-aware generation may limit highly unconventional ideas.
@@
Table 3: Human evaluation of ideas generated by GPT-Researcher with and without metadata ($\%$). P-values are computed with two-tailed Z-tests (* for $p<0.05$, and ** for $p<0.01$).
w. Metadata Tie w/o Metadata P-value
Significance 38.8 22.4 38.8 1.00
Novelty 42.6 14.0 43.4 0.90
Feasibility 46.5 27.1 26.4 0.00**
Exp. Effectiveness 51.2 16.3 32.5 0.00**
Overall 43.4 14.7 41.9 0.80
We perform human evaluation on GPT-Researcher’s output, as this method achieves high rankings in the automatic evaluation. We recruit 18 human annotators at the graduate level or above, with academic backgrounds in social science. For the 50 idea pairs (5 ideas per topic across 10 topics) generated by GPT-Researcher with and without metadata, each pair is annotated by at least two participants. In addition to the four evaluation criteria, we also ask annotators to assess the overall quality of research ideas.
As shown in Table 3, the human evaluation results align with the automatic evaluation trends. The integration of metadata leads to significant improvements in feasibility (20%) and expected effectiveness (18%), though novelty experiences a modest decrease. The overall assessment score increases by 1.5%, suggesting that while metadata strengthens specific quality dimensions, the holistic assessment of research ideas involves balancing multiple criteria. Appendix G demonstrates an example where incorporating metadata improves the feasibility of generated idea, thus improving the overall quality of the idea. We also analyze the agreements between human annotators and LLM judges in Appendix I.
5 Incoporating Automatic Validation in Idea Selection
In this section, we first introduce how we conduct preliminary automatic validation and incorporate the validation results into idea selection, as shown in Figure 1 (bottom). We then assess the effectiveness through: (1) a reference-based evaluation of idea selection performance, and (2) a human evaluation comparing ideas selected with and without the validation process.
@@ 5.1 Incorporating Automatic Validation into Idea Selection
If the idea is deemed testable, the corresponding datasets are then provided to the LLM for validation. We experiment with GPT-4o using the code interpreter assistant, a built-in tool available in GPT models. It achieves superior performance in quantitative reasoning with data [18], while more advanced methods can also be employed in the future. We input the hypotheses along with their corresponding data into the model. The model engages in multi-turn interactions to write and run Python code in a sandbox environment, to validate whether the hypotheses are supported.
We conduct a pilot study to assess the hypothesis validation performance of the approach, and it achieves over 70% accuracy on both general and domain-specific hypotheses. Appendix F provides more details and human evaluation of the validation processes.
@@
We use GPT-4o for both the feasibility check and validation process summarization. Implementation details and prompts are in Appendices E and H. A detailed case of the automatic validation process is in Appendix G.
@@ 5.2 Reference-based Automatic Evaluation
Table 4: Accuracy of judge models in ranking ground-truth ideas among LLM-generated ideas ($\%$). Better results of each judge model are in bold.
@@
We evaluate the impact of validation on idea selection performance, following the setup of ResearchBench [19]. We prompt LLMs to perform pairwise ranking between ground-truth ideas (extracted from academic papers) and LLM-generated ideas, and compare ranking accuracy with and without access to validation processes.
@@
We use Gemini-1.5-Pro and Claude-3.5-Sonnet as judge LLMs. Results are shown in Table 4. For both models, incorporating validation leads to consistently higher average ranking accuracy compared to the setting without validation. Improvements are particularly notable in feasibility and expected effectiveness, while a slight decrease is observed in the judgment of significance and novelty.